# The possible adelic indices for elliptic curves admitting a rational cyclic isogeny

### v1

@@ 4.2. Computation of Possible Images
Proving $7$-adic surjectivity. First, the fiber product $X_{0}(13)\times X_{0}(7)$ (91.112.7.a.1) has no non-CM rational points by Theorem 6. Next, the fiber product $X_{0}(13)\times X_{\mathrm{ns}}^{+}(7)$ has no non-CM rational points by Lemma 17. Lastly, suppose that $E/\mathbb{Q}$ corresponds to a rational point on the fiber product $X_{0}(13)\times X_{\mathrm{sp}}^{+}(7)$. Then there exists a quadratic field $K$ such that $E/K$ corresponds to a quadratic point on the fiber product $X_{\mathrm{sp}}(7)\times X_{0}(13)$. Therefore, $E/K$ corresponds to a quadratic point on the fiber product $X_{0}(7)\times X_{0}(13)$. Moreover, since $E$ is defined over $\mathbb{Q}$ its $j$-invariant lies in $\mathbb{Q}$. From Theorem $1$ of [Vux91], this is impossible.
@@
All of the computations can be verified in the file 13 isogeny l adic images.m in the repository [projectcode]. ∎
@@ 5. A Product Formula for the Adelic Index
Theorem 24 (Theorem 2.7 and Remark 2.8, [MR3350106]).
@@
By [ZywinaOpen, Lemma 1.7], we have that
@@ 6. Handling the Primes $p=11,17$ and $37$
The rational points of $X_{0}(p)$ for $p\in\{11,17,37\}$ are well known in the literature. See page 79 of [MR376533] for $p\in\{11,17\}$ and Section 5, Proposition 2 of [MR354674] for $p=37.$ In this section, we list all of the non-CM $j$-invariants that arise and verify the adelic index for each lies in $\mathcal{I}_{p}$.
@@ 7.1. The Lattice Construction
Our approach follows a lattice-based strategy, drawing inspiration from the articles [MR3500996, MR4468989, ZywinaOpen]. Fix a prime $p\in\{5,7,13\}$ and a subgroup $K\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{p})$ that contains I-I and arises as the $p$-adic Galois image for infinitely many $\overline{\mathbb{Q}}$-isomorphism classes of non-CM elliptic curves over $\mathbb{Q}$ admitting a rational $p$-isogeny. We construct a finite lattice $\mathcal{L}_{p}(K)$, consisting of certain open subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{6})$, recursively as follows:
@@
To enumerate all maximal open subgroups of a given open subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{6})$, we use Corollary 13, which allows us to work at a finite (and computationally tractable) level. It is straightforward to check the conditions in the first bullet point. For the second and third bullet points, we use the results of Section 4: Propositions 21, 22, and 23 for $p=5,7,$ and $13$ respectively.
@@
We compute rational points on the modular curves attached to those groups in $\mathcal{L}_{p}^{\mathrm{fin}}(K)$ that are maximal within this sublattice (i.e., not contained up to conjugacy in any other group of $\mathcal{L}_{p}^{\mathrm{fin}}(K)$). Determining the rational points on these modular curves gives all $6$-adic images containing I-I that occur for only finitely many $\overline{\mathbb{Q}}$-isomorphism classes of non-CM elliptic curves over $\mathbb{Q}$ with $p$-adic image contained in $K$ (up to conjugacy). Such $6$-adic images containing I-I that occur infinitely often are a subset of the groups in $\mathcal{L}_{p}^{\mathrm{inf}}(K)$ (perhaps a proper subset due to the phenomenon of “curious groups” [chiloyan2023classification], but this will not matter for our purposes).
@@ 7.2. Rational Isogenies of Degree 55
Because $K$ arises from an elliptic curve admitting a rational $5$-isogeny, $K$ is necessarily conjugate to a subgroup of 5.6.0.a.1 ($X_{0}(5)$). In Case p = 5.m, we verify that if $\operatorname{SL}_{2}(\mathbb{Z}_{5}):K^{\prime}]=600$, then $K$ is conjugate to a subgroup of either 5.30.0.a.1 ($X_{\mathrm{sp}}(5)$) or 25.30.0.a.1 ($X_{0}(25)$). Thus, to determine the adelic index of any non-CM elliptic curve over $\mathbb{Q}$ admitting a rational $5$-isogeny, it suffices to determine the possible $6$-adic images that can occur along with a $5$-adic image of 5.6.0.a.1 ($X_{0}(5)$), 5.30.0.a.1 ($X_{\mathrm{sp}}(5)$), or 25.30.0.a.1 ($X_{0}(25)$).
We begin by studying the lattice $\mathcal{L}_{5}(X_{0}(5))$ associated with the group 5.6.0.a.1 ($X_{0}(5)$).
@@
In the file Case p = 5.m, we construct the lattice $\mathcal{L}_{5}(X_{0}(5))$ described in Section 7.1 and depicted in Figure 1. The nodes in the diagram represent the groups contained in $\mathcal{L}_{5}(X_{0}(5))$. The white nodes are those in $\mathcal{L}_{5}^{\mathrm{inf}}(X_{0}(5))$, the gray nodes are maximal within $\mathcal{L}_{5}^{\mathrm{fin}}(X_{0}(5))$, and the black nodes are non-maximal within $\mathcal{L}_{5}^{\mathrm{fin}}(X_{0}(5))$. We have that $|\mathcal{L}_{5}^{\mathrm{inf}}(X_{0}(5))|=14$. By computing the commutator index for each of these $14$ groups, we find that
@@
In contrast to the case of $X_{0}(5)$, the lattice associated with $X_{\mathrm{sp}}(5)$ is rather simple:
@@
We construct this lattice in the file Case p = 5.m. We find that $\mathcal{L}_{5}^{\mathrm{inf}}(X_{\mathrm{sp}}(5))=\{\operatorname{GL}_{2}(\mathbb{Z}_{6})\}$. Thus, the only adelic index arising from the sublattice $\mathcal{L}_{5}^{\mathrm{inf}}(X_{\mathrm{sp}}(5))$ by Proposition 27 is
@@
Finally, we consider the lattice associated with $X_{0}(25)$. It has the same tree structure as in the previous case and is shown in Figure 3.
@@
Again, this lattice is constructed in the file Case p = 5.m. Since $\mathcal{L}_{5}^{\mathrm{inf}}(X_{0}(25))=\{\operatorname{GL}_{2}(\mathbb{Z}_{6})\}$, the only adelic index coming from the sublattice $\mathcal{L}_{5}^{\mathrm{inf}}(X_{0}(25))$ is
@@
This completes the proof of Theorem 4 in the case of $p=5$. All of the rational point computations are verified in the file Section 7.2.m in the repository [projectcode].
@@ 7.3. Rational Isogenies of Degree 77
The coloring scheme of the nodes is the same as in the previous section. We have that $|\mathcal{L}_{7}^{\mathrm{inf}}(X_{0}(7))|=9$. By computing the commutator index for each of these groups, we find that
@@
The adelic index of any elliptic curve over $\mathbb{Q}$ with any of these $j$-invariants is $768\in\mathcal{I}_{7}$. The curve 28.64.3.b.1 has genus $3$ and maps to a rank 0 elliptic curve. Using the code ModCrvToEC [JacobJeremycode], we determine that it has exactly eight rational points, of which four are non-CM. Their $j$-invariants are in $\{\tfrac{3^{3}\cdot 13}{2^{2}},\tfrac{-3^{3}\cdot 13\cdot 479^{3}}{16384}\}$ and both give adelic index $768$. The curves 42.48.2.d.1 and 126.48.2.a.1 are both genus $2$ rank 0, so we use the command Chabauty0 in Magma to determine their rational points. The curve 42.48.2.d.1 has $2$ rational points, both of which are cusps. Lastly, the curve 126.48.2.a.1 has no rational points.
This completes the proof of Theorem 4 in the case of $p=7$. All of the rational point computations are verified in the file Section 7.3.m in the repository [projectcode].
@@ 7.4. Rational Isogenies of Degree $13$
$\{[\operatorname{SL}_{2}(\mathbb{Z}_{6}):H^{\prime}]:H\in\mathcal{L}_{13}^{\mathrm{inf}}(X_{0}(13))\}=\{2\}$
@@
It remains to study the modular curves associated with the two gray nodes. These are 78.84.5.b.1 and 312.28.1.a.1. The first curve admits a map to a rank 0 elliptic curve. Using the code ModCrvToEC [JacobJeremycode], we determine that there are exactly two rational points on this curve, both of which are cusps. The second curve, 312.28.1.a.1, also has rank 0 and two rational points, again both of which are cusps.
This completes the proof of Theorem 4 in the case of $p=13$. All of the rational point computations are verified in the file Section 7.4.m in the repository [projectcode].
@@ 8.1. The Product Approach
For $p\in\{2,3\}$, we take a product-based strategy instead of the lattice-based approach from the previous section. Let $E/\mathbb{Q}$ be a non-CM elliptic curve that admits a rational isogeny of degree 2 or 3, under the additional hypothesis that it does not admit any rational isogenies of larger prime degree, consistent with the assumptions in Propositions 19 and 20. By Propositions 19 and 20, we know all possible $2$-adic and $3$-adic Galois images of $E$ that contain I-I. We denote these sets of images by $\mathcal{L}_{2}\coloneqq\mathcal{L}_{2}(p)$ and $\mathcal{L}_{3}\coloneqq\mathcal{L}_{3}(p)$, respectively. By Proposition 27, the adelic index of $E$ depends only on its $6$-adic image. More specifically, we have
@@
where $\delta_{E}$ is an entanglement factor equal to the order of the common quotient group $Q$ of $G_{E}(2^{\infty})^{\prime}$ and $G_{E}(3^{\infty})^{\prime}$, as in Goursat’s lemma.
For each direct product $G=H\times K$ with $H\in\mathcal{L}_{2}$ and $K\in\mathcal{L}_{3}$, we check whether the modular curve $X_{G}$ has finitely many rational points. Among those that do, we take all maximal such groups and determine all of their rational points. We then proceed to the curves with infinitely many rational points. By iterating over all possible $2$-adic images, $3$-adic images, and entanglement factors, one obtains a superset of the possible adelic indices for the elliptic curves under consideration. Finally, we show that any resulting candidate indices lying outside of $\mathcal{I}_{p}$ do not actually occur.
@@ 8.2. Rational Isogenies of Degree 22
We now discuss the computation of the rational points of these $34$ modular curves. All of the claims related to rational points from this section can be verified in the file Section 8.2.m.
First, the following are elliptic curves of rank 0 with only cuspidal or CM rational points:
Next, several of these curves are amenable to Chabauty techniques. The curves 24.36.2.ba.1 and 24.72.2.jo.1 are bielliptic of genus 2 rank 2. They share the same Weierstrass model
$y^{2}=2x^{6}+2.$
From Table 1, this curve has exactly four rational points, all of which are CM. Meanwhile, the curves with labels 24.72.2.hp.1 and 24.72.2.hp.2 are genus 2 rank 0. Using Chabauty0, we compute all of their rational points, finding that all of these are either cusps or CM points.
@@
It has an automorphism $i\colon[x,y,z]\mapsto[-x,-y,z]$. The quotient is a rank 1 elliptic curve. Therefore, there exists an abelian variety of rank 0, which we denote $V$, such that the Jacobian of this curve decomposes as $E\times V$. For any rational point $P$ on 48.72.3.h.1, the point Pi(P)P-i(P) lies in the torsion subgroup of its Jacobian, which we denote by $J_{C}(\mathbb{Q})_{\operatorname{tors}}$. It thus suffices to compute preimages of rational points in $J_{C}(\mathbb{Q})_{\operatorname{tors}}$ under the map $a\colon C(\mathbb{Q})\to J_{C}(\mathbb{Q})_{\operatorname{tors}}$ given by $P\mapsto P-i(P)$. Note that $a$ is injective away from the fixed points of $i$. By using local computations, we can deduce that $J_{C}(\mathbb{Q})_{\operatorname{tors}}$ is a subgroup of $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Further, by working modulo $5$, we check that there is no divisor of the form PQP-Q that has order $2$. Therefore, the only rational points of $C$ are the fixed points of this automorphism. It is easy to check that there are exactly two fixed points: $(1,0,0)$ and (1,1,0)(-1,1,0). One is a cusp and one is a CM point. Next, the curve 48.72.3.i.2 is also of genus 3 and rank 1. It is defined in $\mathbb{P}^{2}$ by the equation
@@
It has an automorphism $i\colon[x,y,z]\mapsto[x,-y,z]$. The quotient is a rank 1 elliptic curve. By a similar analysis, we show that the only rational points of this curve are the fixed points of this automorphism. As before, we find exactly two fixed points (one cusp, one CM). For the curve 48.72.3.i.1(=16.24.0.m.1 $\times$ $X_{\mathrm{ns}}^{+}(3)$), it suffices to compute all the adelic indices of non-CM rational points of 48.72.3.h.1(=16.24.0.l.2 $\times$ $X_{\mathrm{ns}}^{+}(3)$), and we just determined that there are no such points. This suffices because, from Tables 11 and 12 of [MR4868206], if there is a non-CM $E/\mathbb{Q}$ whose $2$-adic image is a subgroup of 16.24.0.m.1, then there exists an $E^{\prime}/\mathbb{Q}$ that is 2-isogenous to $E$ whose $2$-adic image will be a subgroup of 16.24.0.l.2. From Lemma 8, admitting a 2-isogeny does not affect the $3$-adic image. So, if there is a non-CM $E/\mathbb{Q}$ that corresponds to a rational point on 48.72.3.i.1, then there exists an $E^{\prime}/\mathbb{Q}$ that corresponds to a rational point on 48.72.3.h.1. Further, from Corollary 4 of [MR4830946], we know that isogenous elliptic curves share the same adelic index. By a similar argument, for the curve 48.72.3.h.2, it suffices to compute all the adelic indices of non-CM rational points of 48.72.3.i.2, and we have already determined that there are no such points.
The modular curves 16.96.3.ey.1, 16.96.3.ey.2, 16.96.3.fa.1, and 16.96.3.fa.2 are discussed in Section 9.2 of [MR3500996]. Their noncuspidal, non-CM rational points along with their $j$-invariants are given. Each of the four curves has one non-CM $j$-invariant. These are $2^{4}\cdot 17^{3}$, $2^{11}$, $\frac{17^{3}\cdot 241^{3}}{2^{4}}$, and $\frac{257^{3}}{2^{8}}$, respectively. Each of their adelic indices is 384, which lies in the set $\mathcal{I}_{2}$ in Theorem 4. The curves 32.96.3.bh.1 and 32.96.3.bs.1 are discussed in Section 9.4 of [MR3500996], and their noncuspidal, non-CM rational points along with their $j$-invariants are given. Each has one non-CM $j$-invariant. These are $\tfrac{-3^{3}\cdot 5^{3}\cdot 47^{3}\cdot 1217^{3}}{2^{8}\cdot 31^{8}}$ and $\tfrac{3^{3}\cdot 5^{6}\cdot 13^{3}\cdot 23^{3}\cdot 41^{3}}{2^{16}\cdot 31^{4}}$, respectively. The adelic index for these $j$-invariants is also 384.

### v2

@@ 4.2. Computation of Possible Images
Proving $7$-adic surjectivity. First, the fiber product $X_{0}(13)\times X_{0}(7)$ (91.112.7.a.1) has no non-CM rational points by Theorem 5. Next, the fiber product $X_{0}(13)\times X_{\mathrm{ns}}^{+}(7)$ has no non-CM rational points by Lemma 16. Lastly, suppose that $E/\mathbb{Q}$ corresponds to a rational point on the fiber product $X_{0}(13)\times X_{\mathrm{sp}}^{+}(7)$. Then there exists a quadratic field $K$ such that $E/K$ corresponds to a quadratic point on the fiber product $X_{\mathrm{sp}}(7)\times X_{0}(13)$. Therefore, $E/K$ corresponds to a quadratic point on the fiber product $X_{0}(7)\times X_{0}(13)$. Moreover, since $E$ is defined over $\mathbb{Q}$ its $j$-invariant lies in $\mathbb{Q}$. From Theorem $1$ of [43], this is impossible.
@@
All of the computations can be verified in the file 13 isogeny l adic images.m in the repository [16]. ∎
@@ 5. A Product Formula for the Adelic Index
Theorem 23 (Theorem 2.7 and Remark 2.8, [23]).
@@
By [46, Lemma 1.7], we have that
@@ 6. Handling the Primes $p=11,17$ and $37$
The rational points of $X_{0}(p)$ for $p\in\{11,17,37\}$ are well known in the literature. See page 79 of [8] for $p\in\{11,17\}$ and Section 5, Proposition 2 of [33] for $p=37.$ In this section, we list all of the non-CM $j$-invariants that arise and verify the adelic index for each lies in $\mathcal{I}_{p}$.
@@ 7.1. The Lattice Construction
Our approach follows a lattice-based strategy, drawing inspiration from the articles [37, 36, 46]. Fix a prime $p\in\{5,7,13\}$ and a subgroup $K\subseteq\operatorname{GL}_{2}(\mathbb{Z}_{p})$ that contains I-I and arises as the $p$-adic Galois image for infinitely many $\overline{\mathbb{Q}}$-isomorphism classes of non-CM elliptic curves over $\mathbb{Q}$ admitting a rational $p$-isogeny. We construct a finite lattice $\mathcal{L}_{p}(K)$, consisting of certain open subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{6})$, recursively as follows:
@@
To enumerate all maximal open subgroups of a given open subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{6})$, we use Corollary 12, which allows us to work at a finite (and computationally tractable) level. It is straightforward to check the conditions in the first bullet point. For the second and third bullet points, we use the results of Section 4: Propositions 20, 21, and 22 for $p=5,7,$ and $13$ respectively.
@@
We compute rational points on the modular curves attached to those groups in $\mathcal{L}_{p}^{\mathrm{fin}}(K)$ that are maximal within this sublattice (i.e., not contained up to conjugacy in any other group of $\mathcal{L}_{p}^{\mathrm{fin}}(K)$). Determining the rational points on these modular curves gives all $6$-adic images containing I-I that occur for only finitely many $\overline{\mathbb{Q}}$-isomorphism classes of non-CM elliptic curves over $\mathbb{Q}$ with $p$-adic image contained in $K$ (up to conjugacy). Such $6$-adic images containing I-I that occur infinitely often are a subset of the groups in $\mathcal{L}_{p}^{\mathrm{inf}}(K)$ (perhaps a proper subset due to the phenomenon of “curious groups” [12], but this will not matter for our purposes).
@@ 7.2. Rational Isogenies of Degree 55
Because $K$ arises from an elliptic curve admitting a rational $5$-isogeny, $K$ is necessarily conjugate to a subgroup of 5.6.0.a.1 ($X_{0}(5)$). In Case p = 5.m, we verify that if $\operatorname{SL}_{2}(\mathbb{Z}_{5}):K^{\prime}]=600$, then $K$ is conjugate to a subgroup of either 5.30.0.a.1 ($X_{\mathrm{sp}}(5)$) or 25.30.0.a.1 ($X_{0}(25)$). Thus, to determine the adelic index of any non-CM elliptic curve over $\mathbb{Q}$ admitting a rational $5$-isogeny, it suffices to determine the possible $6$-adic images that can occur along with a $5$-adic image of 5.6.0.a.1 ($X_{0}(5)$), 5.30.0.a.1 ($X_{\mathrm{sp}}(5)$), or 25.30.0.a.1 ($X_{0}(25)$). We address each of these now.
7.2.1. The lattice $\mathcal{L}_{5}(X_{0}(5))$
We begin by studying the lattice $\mathcal{L}_{5}(X_{0}(5))$ associated with the group 5.6.0.a.1 ($X_{0}(5)$). In the file Case p = 5.m, we construct the lattice $\mathcal{L}_{5}(X_{0}(5))$ described in Section 7.1 and depicted in Figure 1. The nodes in the diagram represent the groups contained in $\mathcal{L}_{5}(X_{0}(5))$. The white nodes are those in $\mathcal{L}_{5}^{\mathrm{inf}}(X_{0}(5))$, the gray nodes are maximal within $\mathcal{L}_{5}^{\mathrm{fin}}(X_{0}(5))$, and the black nodes are non-maximal within $\mathcal{L}_{5}^{\mathrm{fin}}(X_{0}(5))$.
@@ 7.2.1. The lattice $\mathcal{L}_{5}(X_{0}(5))$
We begin by considering the sublattice $\mathcal{L}_{5}^{inf}(X_{0}(5))$. We have that $|\mathcal{L}_{5}^{\mathrm{inf}}(X_{0}(5))|=14$. By computing the commutator index for each of these $14$ groups, we find that
@@
7.2.2. The lattice $\mathcal{L}_{5}(X_{sp}(5))$
In contrast to the case of $X_{0}(5)$, the lattice associated with $X_{\mathrm{sp}}(5)$ is rather simple. We construct this lattice in the file Case p = 5.m.
@@ 7.2.2. The lattice $\mathcal{L}_{5}(X_{sp}(5))$
Again, we begin by considering the one node that constitutes $\mathcal{L}_{5}^{\mathrm{inf}}$. We find that $\mathcal{L}_{5}^{\mathrm{inf}}(X_{\mathrm{sp}}(5))=\{\operatorname{GL}_{2}(\mathbb{Z}_{6})\}$. Thus, the only adelic index arising from the sublattice $\mathcal{L}_{5}^{\mathrm{inf}}(X_{\mathrm{sp}}(5))$ by Proposition 26 is
@@
7.2.3. The lattice $\mathcal{L}_{5}(X_{0}(25))$
Finally, we consider the lattice associated with $X_{0}(25)$. It has the same tree structure as in the previous case and is shown in Figure 3. Again, this lattice is constructed in the file Case p = 5.m.
@@ 7.2.3. The lattice $\mathcal{L}_{5}(X_{0}(25))$
Since $\mathcal{L}_{5}^{\mathrm{inf}}(X_{0}(25))=\{\operatorname{GL}_{2}(\mathbb{Z}_{6})\}$, the only adelic index coming from the sublattice $\mathcal{L}_{5}^{\mathrm{inf}}(X_{0}(25))$ is
@@
This completes the proof of Theorem 4 in the case of $p=5$. All of the rational point computations are verified in the file Section 7.2.m in the repository [16].
@@ 7.3. Rational Isogenies of Degree 77
The coloring scheme (perhaps it would be more accurate to say “shading scheme”) of the nodes is the same as in the previous section. We have that $|\mathcal{L}_{7}^{\mathrm{inf}}(X_{0}(7))|=9$. By computing the commutator index for each of these groups, we find that
@@
The adelic index of any elliptic curve over $\mathbb{Q}$ with any of these $j$-invariants is $768\in\mathcal{I}_{7}$. The curve 28.64.3.b.1 has genus $3$ and maps to a rank 0 elliptic curve. Using the code ModCrvToEC [31], we determine that it has exactly eight rational points, of which four are non-CM. Their $j$-invariants are in $\{\tfrac{3^{3}\cdot 13}{2^{2}},\tfrac{-3^{3}\cdot 13\cdot 479^{3}}{16384}\}$ and both give adelic index $768$. The curves 42.48.2.d.1 and 126.48.2.a.1 are both genus $2$ rank 0, so we use the command Chabauty0 in Magma to determine their rational points. The curve 42.48.2.d.1 has $2$ rational points, both of which are cusps. Lastly, the curve 126.48.2.a.1 has no rational points.
This completes the proof of Theorem 4 in the case of $p=7$. All of the rational point computations are verified in the file Section 7.3.m in the repository [16].
@@ 7.4. Rational Isogenies of Degree $13$
$\{[\operatorname{SL}_{2}(\mathbb{Z}_{6}):H^{\prime}]:H\in\mathcal{L}_{13}^{\mathrm{inf}}(X_{0}(13))\}=\{2\}.$
@@
It remains to study the modular curves associated with the two gray nodes. These are 78.84.5.b.1 and 312.28.1.a.1. The first curve admits a map to a rank 0 elliptic curve. Using the code ModCrvToEC [31], we determine that there are exactly two rational points on this curve, both of which are cusps. The second curve, 312.28.1.a.1, also has rank 0 and two rational points, again both of which are cusps.
This completes the proof of Theorem 4 in the case of $p=13$. All of the rational point computations are verified in the file Section 7.4.m in the repository [16].
@@ 8.1. The Product Approach
For $p\in\{2,3\}$, we take a product-based strategy instead of the lattice-based approach from the previous section. Let $E/\mathbb{Q}$ be a non-CM elliptic curve that admits a rational isogeny of degree 2 or 3, under the additional hypothesis that it does not admit any rational isogenies of larger prime degree, consistent with the assumptions in Propositions 18 and 19. By Propositions 18 and 19, we know all possible $2$-adic and $3$-adic Galois images of $E$ that contain I-I. We denote these sets of images by $\mathcal{L}_{2}\coloneqq\mathcal{L}_{2}(p)$ and $\mathcal{L}_{3}\coloneqq\mathcal{L}_{3}(p)$, respectively. Assuming the $\ell$-adic Galois representation is surjective for all $\ell\geq 5$, by Proposition 26, the adelic index of $E$ depends only on its $6$-adic image. More specifically, in this case we have
@@
where $\delta_{E}$ is the entanglement factor, defined as the order of the common quotient group $Q$ of $G_{E}(2^{\infty})^{\prime}$ and $G_{E}(3^{\infty})^{\prime}$, as in Goursat’s lemma. Proposition 18 guarantees that the $\ell$-adic representation is surjective for all $\ell\geq 5$ in the $p=2$ case; in the $p=3$ case, it is possible that the $5$-adic representation is nonsurjective, and we deal with this possibility at the beginning of Section 8.3.
For each direct product $G=H\times K$ with $H\in\mathcal{L}_{2}$ and $K\in\mathcal{L}_{3}$, we check whether the modular curve $X_{G}$ has finitely many rational points. Among those that do, we take all maximal such groups and determine all of their rational points (or explain why doing so is unnecessary for our purposes). We then proceed to the curves with infinitely many rational points. By iterating over all possible $2$-adic images, $3$-adic images, and entanglement factors, one obtains a superset of the possible adelic indices for the elliptic curves under consideration. Finally, we show that any resulting candidate indices lying outside of $\mathcal{I}_{p}$ do not actually occur.
@@ 8.2. Rational Isogenies of Degree 22
8.2.1. Maximal Products with Finitely Many Points
@@
Most of the rational points computations for these curves are straightforward using techniques already discussed earlier in this paper. For these, we provide full details in the file Section 8.2.m. We now discuss the computation for the curves that require more ad hoc techniques.
@@
It has an automorphism $i\colon[x:y:z]\mapsto[-x:-y:z]$. The quotient is a rank 1 elliptic curve. Therefore, there exists an abelian variety of rank 0, which we denote $V$, such that the Jacobian of this curve decomposes as $E\times V$. For any rational point $P$ on 48.72.3.h.1, the point Pi(P)P-i(P) lies in the torsion subgroup of its Jacobian, which we denote by $J_{C}(\mathbb{Q})_{\operatorname{tors}}$. It thus suffices to compute preimages of rational points in $J_{C}(\mathbb{Q})_{\operatorname{tors}}$ under the map $a\colon C(\mathbb{Q})\to J_{C}(\mathbb{Q})_{\operatorname{tors}}$ given by $P\mapsto P-i(P)$. Note that $a$ is injective away from the fixed points of $i$. By using local computations, we can deduce that $J_{C}(\mathbb{Q})_{\operatorname{tors}}$ is a subgroup of $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Further, by working modulo $5$, we check that there is no divisor of the form PQP-Q that has order $2$. Therefore, the only rational points of $C$ are the fixed points of this automorphism. It is easy to check that there are exactly two fixed points: $(1:0:0)$ and (1:1:0)(-1:1:0). One is a cusp and one is a CM point. Next, the curve 48.72.3.i.2 is also of genus 3 and rank 1. It is defined in $\mathbb{P}^{2}$ by the equation
@@
It has an automorphism $i\colon(x:y:z)\mapsto(x:-y:z)$. The quotient is a rank 1 elliptic curve. By a similar analysis, we show that the only rational points of this curve are the fixed points of this automorphism. As before, we find exactly two fixed points (one cusp, one CM). For the curve 48.72.3.i.1(=16.24.0.m.1 $\times$ $X_{\mathrm{ns}}^{+}(3)$), it suffices to compute all the adelic indices of non-CM rational points of 48.72.3.h.1(=16.24.0.l.2 $\times$ $X_{\mathrm{ns}}^{+}(3)$), and we just determined that there are no such points. This suffices because, from Tables 11 and 12 of [13], if there is a non-CM $E/\mathbb{Q}$ whose $2$-adic image is a subgroup of 16.24.0.m.1, then there exists an $E^{\prime}/\mathbb{Q}$ that is 2-isogenous to $E$ whose $2$-adic image will be a subgroup of 16.24.0.l.2. From Lemma 7, admitting a 2-isogeny does not affect the $3$-adic image. So, if there is a non-CM $E/\mathbb{Q}$ that corresponds to a rational point on 48.72.3.i.1, then there exists an $E^{\prime}/\mathbb{Q}$ that corresponds to a rational point on 48.72.3.h.1. Further, from Corollary 4 of [21], we know that isogenous elliptic curves share the same adelic index. By a similar argument, for the curve 48.72.3.h.2, it suffices to compute all the adelic indices of non-CM rational points of 48.72.3.i.2, and we have already determined that there are no such points.
The modular curves 16.96.3.ey.1, 16.96.3.ey.2, 16.96.3.fa.1, and 16.96.3.fa.2 are discussed in Section 9.2 of [37]. Their noncuspidal, non-CM rational points along with their $j$-invariants are given. Each of the four curves has one non-CM $j$-invariant. These are $2^{4}\cdot 17^{3}$, $2^{11}$, $\frac{17^{3}\cdot 241^{3}}{2^{4}}$, and $\frac{257^{3}}{2^{8}}$, respectively. Each of their adelic indices is 384, which lies in the set $\mathcal{I}_{2}$ in Theorem 4. The curves 32.96.3.bh.1 and 32.96.3.bs.1 are discussed in Section 9.4 of [37], and their noncuspidal, non-CM rational points along with their $j$-invariants are given. Each has one non-CM $j$-invariant. These are $\tfrac{-3^{3}\cdot 5^{3}\cdot 47^{3}\cdot 1217^{3}}{2^{8}\cdot 31^{8}}$ and $\tfrac{3^{3}\cdot 5^{6}\cdot 13^{3}\cdot 23^{3}\cdot 41^{3}}{2^{16}\cdot 31^{4}}$, respectively. The adelic index for these $j$-invariants is also 384.
@@
8.2.2. Products with Infinitely Many Rational Points